# Comparative study of the butterfly velocity in holographic QCD models at finite temperature and chemical potential

### v2

@@ 3.0.1 Background
Both 1RCBH and 2RCBH models are cases of a more general top-down holographic model called the STU model [82, 83]. The STU gravity model describes five-dimensional black brane solutions containing three conserved Abelian $R$-charges. One gets the 1RCBH model when two $R$-charges are set to zero, whereas the 2RCBH model is obtained by setting one $R$-charge to zero and further identifying the remaining two $R$-charges. Therefore, in each case, there is only one nontrivial charge and hence the chemical potential. The black hole solutions of these two models (discussed below) correspond to strongly coupled charged Supersymmetric Yang-Mills (SYM) plasma in the dual boundary side.
@@
where subscripts $1$ and $2$ are used for 1RCBH and 2RCBH models, respectively. Here $L$ is the AdS radius, which we will set to one from now on for convenience. For the following Ansätze for the metric, gauge field, and dilaton field
@@
In the above solutions, $z_{h}$ is the black hole horizon, determined by the largest real root of the blackening function $H_{i}(z)$, and $Q_{i}$ are the black hole charges. The chemical potential of the dual boundary theory $\mu_{i}$, corresponding to the leading order term of the zeroth component of the gauge field at the asymptotic boundary, is related to the black hole charge in the following way
@@
where $V_{3}$ is the unit volume of the three-dimensional boundary space. Note that in the above solutions, $z$ corresponds to the usual holographic radial coordinate and it runs from $z=z_{h}$ (horizon radius) to $z=\infty$ (asymptotic boundary).
Using the above thermodynamic observables, one can analyze the thermodynamic behavior and stability of 1RCBH and 2RCBH black brane solutions. It turns out that there are two black hole branches – large and small – in 1RCBH solution. The large black hole branch (large $z_{h}$ solution) is thermodynamically stable and exhibits a positive specific heat, whereas the small black hole branch (large $z_{h}$ solution) is thermodynamically unstable and exhibits negative specific heat. There also appears a critical point $\mu_{1}=\pi T_{1}/\sqrt{2}$ at which both branches merge and lead to diverging specific heat. Accordingly, the range of values of $\mu_{1}/T_{1}$ probed by the 1RCBH model is restricted to $\mu_{1}/T_{1}\in[0,\pi/\sqrt{2}]$. On the other hand, there is only one black hole branch in the 2RCBH model, which is always thermodynamics stable. Moreover, the allowed range of $\mu_{2}/T_{2}$ is unrestricted in the 2RCBH model, i.e., it probes values of $\mu_{2}/T_{2}$ ranging from zero to infinity. Below, in the computation of the butterfly velocity, we will work in the thermodynamically stable phases and restrict ourselves to only those values of the chemical potential and temperature that are allowed in 1RCBH and 2RCBH black brane solutions.
3.0.2 Calculation of $v_{B}$ using the entanglement wedge method
We first compute the butterfly velocity for the 1RCBH model. Here, we concentrate only on the large black hole solution, corresponding to a stable thermal state. Comparing the 1RCBH metric (3.0.1) with the general metric (2.1), we have the following for the 1RCBH model: [^3] Here $f(z)$ should not be confused with the coupling functions $\tilde{f}_{1}(\phi)$ and $\tilde{f}_{2}(\phi)$.
@@ 3.0.2 Calculation of $v_{B}$ using the entanglement wedge method
Computing the metric functions and their derivatives at the horizon and substituting in (2.13), we get the following expression for the butterfly velocity
@@
We can express $v_{B}$ in terms of $\mu_{1}$ by substituting $Q_{1}$ from (3.6) into (3.9)
@@
Substituting (3.11) into (3.10), we get the following expression of $v_{B}$ in terms of $\mu_{1}$ and $T_{1}$ for the 1RCBH model
@@
Substituting the near-horizon values of these metric functions into (2.13) and further expressing $Q_{2}$ in terms of $\mu_{2}$ from (3.6), we get the following expression for the butterfly velocity for the 2RCBH model
@@
In Figure 3, we illustrate the chemical potential and thermal dependent profiles of the butterfly velocity in the 1RCBH model. We find that $v_{B}^{2}$ monotonically decreases with $\mu_{1}$ for all temperatures. This suggests that the rate at which information about the perturbation propagates among the local degrees of freedom in the dual boundary theory decreases as one increases the chemical potential. In Figure 3, we have only plotted the thermodynamically stable solution. Note that the range of values of $\mu_{1}/T_{1}$ probed by the 1RCBH model is $[0,\pi/\sqrt{2}]$ and this is automatically taken care of, due to the presence of the square root ($\sqrt{\pi^{2}T_{1}^{2}-2\mu_{1}^{2}}$) in (3.11), as is also evident from the figures. We find that in the allowed range of parameters $\mu_{1}$ and $T_{1}$, $v_{B}^{2}$ always exhibits a monotonically decreasing profile with respect to $\mu_{1}$.
@@
Similarly, $v_{B}^{2}$ is found to increase monotonically with $T_{1}$ for all chemical potential values. In particular, $v_{B}^{2}$ increases rapidly at low temperatures and then saturates to a constant value at high temperatures. This is an expected result, as one can expect that the local degrees of freedom would diffuse faster as one increases the temperature. It is interesting to note that $v_{B}$ asymptotes to $v_{B}^{S}$, with $v_{B}^{S}$ being the butterfly velocity in the background of the AdS-Schwarzschild black brane, in the high temperature limit ($T_{1}\rightarrow\infty$) for all values of $\mu$, i.e.,
@@
Note that in the case of chargeless plasma, by considering the AdS-Schwarzschild black brane, $v_{B}^{2}$ was found to be independent of temperature and takes a constant value [67]. With the chemical potential, its thermal structure becomes much more richer, as is evident from Figure 3. For relativistic theories, $v_{B}^{S}$ gives the upper bound on the butterfly velocity [67]. It is interesting to see that this bound is always respected and only gets saturated in the high temperature limit in the 1RCBH model. This might be because at very high temperatures, all other scale gets suppressed, leading to a scale symmetry. As we will see in the following sections, this result remains true in other holographic plasmas as well.
The profile of the butterfly velocity in the 2RCBH model is shown in Figure 4. Most of the results of $v_{B}^{2}$ here are similar to the 1RCBH model. In particular, again $v_{B}^{2}$ decreases monotonically with the chemical potential at all fixed temperatures. Similarly, $v_{B}^{2}$ increases rapidly at low temperatures and then saturates to a constant value at high temperatures. Just like in the case of the 1RCBH model, $v_{B}$ for the 2RCBH model again asymptotes to $v_{B}^{S}$ in the high temperature limit. However, there are also some differences. For instance, in contrast to the 1RCBH model, there are no restrictions on the allowed values of temperature and chemical potential in the 2RCBH model. Accordingly, one can probe $v_{B}^{2}$ in a much larger parameter space of $(T_{2},~{}\mu_{2})$ in 2RCBH model. This leads to some differences between the 1RCBH and 2RCBH models. In particular, at any fixed temperature, $v_{B}^{2}$ asymptotes to zero as the chemical potential is taken to infinity. This can be explicitly observed from Eq. (3.14). However, this was not the case in the 1RCBH model, where $v_{B}^{2}$ attained a temperature dependent finite minimum value. Similarly, at a fixed finite chemical potential, $v_{B}^{2}$ is bounded from below by a positive number in the 1RCBH model, whereas in the 2RCBH model, $v_{B}^{2}$ can be vanishingly small. Moreover, we further find that for the same values of temperature and chemical potential, $v_{B}^{2}$ is higher in the 2RCBH model compared to the 1RCBH model, suggesting information propagates faster in the former compared to the latter model.
@@
3.0.3 Calculation of $v_{B}$ using OTOC method
The metric functions for the 1RCBH model are given in (LABEL:eqn:metricfunctions1RCBH). Substituting these metric functions into (2.22), and using (2.25), we get the following expressions for $A(u,v)=A_{1}(u,v)$ and $B(u,v)=B_{1}(u,v)$ in the 1RCBH model
@@ 3.0.3 Calculation of $v_{B}$ using OTOC method
Computing the limits of $A_{1}$ and $\partial_{u}\partial_{v}B_{1}$ as $u\rightarrow 0$, we get the following
@@
with $\alpha_{1}=2\pi T_{1}$. Finally, putting everything into (2.35), we get the following expression of $v_{B}$
@@
where we have traded $Q_{1}$ for $\mu_{1}$. The above expression matches with the analogous expression obtained for $v_{B}^{2}$ from the entanglement wedge method in the previous section (3.10).
@@
3.0.4 Calculation of $v_{B}$ using pole-skipping method
@@
Now, we differentiate (3.29) to obtain $\delta T_{\alpha\beta}$
$\begin{split}\delta T_{\alpha\beta}=&\mathcal{L}_{M}\delta g_{\alpha\beta}+g_{\alpha\beta}\left(\frac{\delta\mathcal{L}_{M}}{\delta g^{\rho\lambda}}\delta g^{\rho\lambda}+\frac{\delta\mathcal{L}_{M}}{\delta\phi}\delta\phi+\frac{\delta\mathcal{L}_{M}}{\delta A^{\rho}}\delta A^{\rho}\right)+(\partial_{\alpha}\delta\phi)(\partial_{\beta}\phi)\\ &+(\partial_{\alpha}\phi)(\partial_{\beta}\delta\phi)+\tilde{f}^{\prime}(\phi)F_{\alpha\rho}F_{\beta\lambda}g^{\rho\lambda}\delta\phi-\tilde{f}(\phi)F_{\alpha\mu}F_{\beta\sigma}g^{\mu\rho}g^{\lambda\sigma}\delta g_{\rho\lambda}\\ &+\tilde{f}(\phi)g^{\rho\sigma}\left(F_{\alpha\rho}(\partial_{\beta}\delta A_{\sigma}-\partial_{\sigma}\delta A_{\beta})+(\partial_{\alpha}\delta A_{\rho}-\partial_{\rho}\delta A_{\alpha})F_{\beta\sigma}\right)\,,\end{split}$ (3.31)
@@
Here, we have used the fact that $g^{zz}(z_{h})=g_{vv}(z_{h})=0$. Substituting Eqs. (3.30) and (3.32) into (2.44) and using the fact that $g^{vz}(z_{h})=1/\sqrt{h(z_{h})f(z_{h})}$ and $g_{vz}(z_{h})=\sqrt{h(z_{h})f(z_{h})}$, it is straightforward to show that the equation (2.44) holds true for both 1RCBH and 2RCBH models.
For completeness, we also compute the pole-skipping points of 1RCBH and 2RCBH models. Substituting metric functions from (LABEL:eqn:metricfunctions1RCBH) into (2.47), we get the following pole-skipping points in the 1RCBH model
@@
From the above, we get a butterfly velocity that matches the results of the other two methods:
@@ 4.1 Background
In this section, we turn our attention to a bottom-up holographic EMD model and investigate the butterfly velocity at finite temperatures and chemical potentials in the dual QCD deconfined plasma. As mentioned in the introduction, bottom-up models generally mimic QCD properties more accurately, as compared to their higher-dimensional string theory-inspired top-down counterparts, and therefore, it is interesting to analyze the butterfly velocity in bottom-up models as well. To make our analysis more complete, we will consider another sophisticated bottom-up EMD based QCD model in the next section.
Here we consider the bottom-up holographic QCD model of [52, 53]. The most general version of this model is a solution to the EMD gravity (3.1) containing Maxwell and dilaton matter fields. This model is based on the potential reconstruction technique [84, 85, 86, 87, 88, 89, 90, 91], where the potential is self-consistently fixed from field equations. In particular, for the following Ansätze for the metric, dilaton, and gauge fields:
$ds^{2}=e^{2A(z)}\left(-z^{2}G(z)dt^{2}+\frac{dz^{2}}{z^{2}G(z)}+z^{2}(dx_{1}^{2}+dx^{2}+dx_{3}^{2})\right)\,,$ (4.1)
$\phi=\phi(z)\,,~{}~{}~{}\mathcal{A}_{M}=\Phi(z)\delta_{M}^{0}\,,$
@@
Here $\mu$ is the chemical potential, $a$ and $c$ are model parameters that are fixed by taking inputs from real QCD. In particular, by demanding the deconfinement transition to be around $0.270~{}GeV$, as in the pure glue sector, fixes $a=0.15~{}\text{GeV}^{2}$, while the requirement of the linear Regge trajectory in the heavy-meson spectrum fixes $c=1.16~{}\text{GeV}^{2}$. Again, $z$ is the usual radial coordinate that runs from $z=z_{h}$ (horizon) to $z=\infty$ (boundary). The entropy and temperature of the black hole are given by:
@@
There is another solution to this gravity model. This second solution corresponds to thermal-AdS, which does not have a horizon. The thermodynamic phase structure of the above solution reveals that the black hole solution exists only above a minimum temperature $T_{min}$, and below $T_{min}$, the black hole solution ceases to exist. Interestingly, there exists a Hawking/Page type phase transition between thermal-AdS and black hole solutions as the temperature is varied. In particular, the black hole phase is thermodynamically favored at high temperatures, whereas the thermal-AdS phase is favored at low temperatures. The corresponding phase transition temperature $T_{c}$ is a chemical potential dependent quantity, and in particular, it decreases as the chemical potential increases. In the dual boundary theory, the thermal-AdS and black hole phases correspond to the confined and deconfined phases. Accordingly, this EMD gravity model provides a self-consistent bottom-up holographic model for the deconfinement transition, which mimics real QCD behavior quite well. Since our main aim is to analyze the butterfly velocity and its thermal and chemical potential dependent profile in the deconfined plasma phase, we will mainly concentrate on the black hole solution in the next subsections.

### v3

@@ 3.1 Background
Both 1RCBH and 2RCBH models are cases of a more general top-down holographic model called the STU model [83, 84]. The STU gravity model describes five-dimensional black brane solutions containing three conserved Abelian $R$ charges. The 1RCBH model is obtained when two $R$ charges are set to zero, whereas the 2RCBH model is obtained by setting one $R$ charge to zero and further identifying the remaining two $R$ charges. Therefore, in each case, there is only one nontrivial charge and hence the chemical potential. Essentially, the dual field theory configurations of the 1RCBH model should be thought of as states in $\mathcal{N}=4$ super-Yang-Mills (SYM) theory with a temperature and a chemical potential for a $U(1)$ subgroup of the $SO(6)$ $R$ symmetry. Similarly, the 2RCBH model corresponds to a two-charge $\mathcal{N}=4$ solution, which has equal charges for two $U(1)$ gauge fields inside the $SO(6)$ $R$ symmetry. The black hole solutions of these two models (discussed below) correspond to a strongly coupled charged SYM plasma in the dual boundary side.
The two models represent different realizations of the strongly coupled $\mathcal{N}=4$ SYM plasma at finite temperature and charge density. Since these models are analytically solvable, and geometric and thermodynamic properties are known analytically, they have been greatly used in holographic literature. Both exhibit several features in common with QCD-like black holes, i.e., models where the dilaton potential and gauge-kinetic coupling are tuned to reproduce lattice QCD results [51]. The 1RCBH model features a critical point and has been used to study the QCD phase diagram and thermodynamics [51]. Various transport phenomena relevant to QCD, such as conductivity and diffusion, have also been investigated within this model [51]. In addition to thermodynamics and transport, several quantum information observables [85] [86] and chaotic properties [87] [66] have been explored. In this work, we provide a more comprehensive analysis of the model’s chaotic behavior. On the other hand, the 2RCBH model does not exhibit any critical point. Nevertheless, it has been employed to study a range of transport phenomena in QCD [52], nonequilibrium homogeneous isotropization dynamics [88], nonhydrodynamic quasinormal modes and late-time equilibration of the plasma [89]. Moreover, this model captures non-Fermi liquid behavior, exhibiting nested Fermi surfaces with non-quasi-particle excitations, phenomena which could be potentially relevant to dense QCD matter [52]. To the best of our knowledge, the chaotic properties of the 2RCBH model have not been previously explored. Our work is, therefore, the first to present a detailed analysis of chaos in the 2RCBH model.
@@
where subscripts $1$ and $2$ are used for 1RCBH and 2RCBH models, respectively. Here $L$ is the AdS radius, which we will set to $1$ from now on for convenience. For the following Ansätze for the metric, gauge field, and dilaton field:
@@
In the above solutions, $z_{h}$ is the black hole horizon, determined by the largest real root of the blackening function $H_{i}(z)$, and $Q_{i}$ are the black hole charges. The chemical potential of the dual boundary theory $\mu_{i}$, corresponding to the leading-order term of the zeroth component of the gauge field at the asymptotic boundary, is related to the black hole charge in the following way:
@@
where $V_{3}$ is the unit volume of the three-dimensional boundary space. Note that, in the above solutions, $z$ corresponds to the usual holographic radial coordinate and it runs from $z=z_{h}$ (horizon radius) to $z=\infty$ (asymptotic boundary).
Using the above thermodynamic observables, one can analyze the thermodynamic behavior and stability of 1RCBH and 2RCBH black brane solutions. It turns out that there are two black hole branches - large and small - in the 1RCBH solution. The large black hole branch (large $z_{h}$ solution) is thermodynamically stable and exhibits a positive specific heat, whereas the small black hole branch (large $z_{h}$ solution) is thermodynamically unstable and exhibits negative specific heat. There also appears a critical point $\mu_{1}=\pi T_{1}/\sqrt{2}$ at which both branches merge and lead to diverging specific heat. Accordingly, the range of values of $\mu_{1}/T_{1}$ probed by the 1RCBH model is restricted to $\mu_{1}/T_{1}\in[0,\pi/\sqrt{2}]$. As discussed in [90], $\mu/T=\pi/\sqrt{2}$ is a critical point of second-order phase transition. Note that this phase transition is not the usual Hawking/Page phase transition. On the other hand, there is only one black hole branch in the 2RCBH model, which is always thermodynamically stable. Moreover, the allowed range of $\mu_{2}/T_{2}$ is unrestricted in the 2RCBH model; i.e., it probes values of $\mu_{2}/T_{2}$ ranging from zero to infinity. Below, in the computation of the butterfly velocity, we will work in the thermodynamically stable phases and restrict ourselves to only those values of the chemical potential and temperature that are allowed in 1RCBH and 2RCBH black brane solutions.
3.2 Calculation of $v_{B}$ using the entanglement wedge method
We first compute the butterfly velocity for the 1RCBH model. Here, we concentrate only on the large black hole solution, corresponding to a stable thermal state. Comparing the 1RCBH metric (3.1) with the general metric (2.1), we have the following for the 1RCBH model [^3] Here $f(z)$ should not be confused with the coupling functions $\tilde{f}_{1}(\phi)$ and $\tilde{f}_{2}(\phi)$.:
@@ 3.2 Calculation of $v_{B}$ using the entanglement wedge method
Computing the metric functions and their derivatives at the horizon and substituting in (2.13), we get the following expression for the butterfly velocity:
@@
We can express $v_{B}$ in terms of $\mu_{1}$ by substituting $Q_{1}$ from (3.6) into (3.9),
@@
Substituting (3.11) into (3.10), we get the following expression of $v_{B}$ in terms of $\mu_{1}$ and $T_{1}$ for the 1RCBH model:
@@
Substituting the near-horizon values of these metric functions into (2.13) and further expressing $Q_{2}$ in terms of $\mu_{2}$ from (3.6), we get the following expression for the butterfly velocity for the 2RCBH model:
@@
In Figure 3, we illustrate the chemical potential and temperature-dependent profiles of the butterfly velocity in the 1RCBH model. We find that $v_{B}^{2}$ monotonically decreases with $\mu_{1}$ for all temperatures. This suggests that the rate at which information about the perturbation propagates among the local degrees of freedom in the dual boundary theory decreases as one increases the chemical potential. In Figure 3, we have only plotted the thermodynamically stable solution. Note that the range of values of $\mu_{1}/T_{1}$ probed by the 1RCBH model is $0,\pi/\sqrt{2}]$ and this is automatically taken care of, due to the presence of the square root ($\sqrt{\pi^{2}T_{1}^{2}-2\mu_{1}^{2}}$) in (3.11), as is also evident from the figures. We find that, in the allowed range of parameters $\mu_{1}$ and $T_{1}$, $v_{B}^{2}$ always exhibits a monotonically decreasing profile with respect to $\mu_{1}$.
@@
Similarly, $v_{B}^{2}$ is found to increase monotonically with $T_{1}$ for all chemical potential values. In particular, $v_{B}^{2}$ increases rapidly at low temperatures and then saturates to a constant value at high temperatures. This is an expected result, as one can expect that the local degrees of freedom would diffuse faster as one increases the temperature. It is interesting to note that $v_{B}$ asymptotes to $v_{B}^{S}$, with $v_{B}^{S}$ being the butterfly velocity in the background of the AdS-Schwarzschild black brane, in the high-temperature limit ($T_{1}\rightarrow\infty$) for all values of $\mu$, i.e.,
@@
Note that, in the case of chargeless plasma, by considering the AdS-Schwarzschild black brane, $v_{B}^{2}$ was found to be independent of temperature and takes a constant value [68]. With the chemical potential, its thermal structure becomes much richer, as is evident from Figure 3. For relativistic theories, $v_{B}^{S}$ gives the upper bound on the butterfly velocity [68]. It is interesting to see that this bound is always respected and only gets saturated in the high-temperature limit in the 1RCBH model. This might be because, at very high temperatures, all other scales get suppressed, leading to a scale symmetry. As we will see in the following sections, this result remains true in other holographic plasmas as well.
The profile of the butterfly velocity in the 2RCBH model is shown in Figure 4. Most of the results of $v_{B}^{2}$ here are similar to the 1RCBH model. In particular, again $v_{B}^{2}$ decreases monotonically with the chemical potential at all fixed temperatures. Similarly, $v_{B}^{2}$ increases rapidly at low temperatures and then saturates to a constant value at high temperatures. Just like in the case of the 1RCBH model, $v_{B}$ for the 2RCBH model again asymptotes to $v_{B}^{S}$ in the high-temperature limit. However, there are also some differences. For instance, in contrast to the 1RCBH model, there are no restrictions on the allowed values of temperature and chemical potential in the 2RCBH model. Accordingly, one can probe $v_{B}^{2}$ in a much larger parameter space of $(T_{2},~\mu_{2})$ in 2RCBH model. This leads to some differences between the 1RCBH and 2RCBH models. In particular, at any fixed temperature, $v_{B}^{2}$ asymptotes to zero as the chemical potential is taken to infinity. This can be explicitly observed from Eq. (3.14). However, this was not the case in the 1RCBH model, where $v_{B}^{2}$ attained a temperature-dependent finite minimum value. Similarly, at a fixed finite chemical potential, $v_{B}^{2}$ is bounded from below by a positive number in the 1RCBH model, whereas in the 2RCBH model, $v_{B}^{2}$ can be vanishingly small. Moreover, we further find that, for the same values of temperature and chemical potential, $v_{B}^{2}$ is higher in the 2RCBH model compared to the 1RCBH model, suggesting information propagates faster in the former compared to the latter model.
@@
3.3 Calculation of $v_{B}$ using OTOC method
The metric functions for the 1RCBH model are given in (LABEL:eqn:metricfunctions1RCBH). Substituting these metric functions into (2.22) and using (2.25), we get the following expressions for $A(u,v)=A_{1}(u,v)$ and $B(u,v)=B_{1}(u,v)$ in the 1RCBH model:
@@ 3.3 Calculation of $v_{B}$ using OTOC method
Computing the limits of $A_{1}$ and $\partial_{u}\partial_{v}B_{1}$ as $u\rightarrow 0$, we get the following:
@@
with $\alpha_{1}=2\pi T_{1}$. Finally, putting everything into (2.35), we get the following expression of $v_{B}$:
@@
where we have traded $Q_{1}$ for $\mu_{1}$. The above expression matches with the analogous expression obtained for $v_{B}^{2}$ from the entanglement wedge method in the previous section [Eq. (3.10)].
@@
3.4 Calculation of $v_{B}$ using pole-skipping method
@@
Now, we differentiate (3.29) to obtain $\delta T_{\alpha\beta}$,
[…]
@@
Here, we have used the fact that $g^{zz}(z_{h})=g_{vv}(z_{h})=0$. Substituting Eqs. (3.30) and (3.32) into (2.44) and using the fact that $g^{vz}(z_{h})=1/\sqrt{h(z_{h})f(z_{h})}$ and $g_{vz}(z_{h})=\sqrt{h(z_{h})f(z_{h})}$, it is straightforward to show that Eq. (2.44) holds true for both 1RCBH and 2RCBH models.
For completeness, we also compute the pole-skipping points of 1RCBH and 2RCBH models. Substituting metric functions from (LABEL:eqn:metricfunctions1RCBH) into (2.47), we get the following pole-skipping points in the 1RCBH model:
@@
From the above, we get a butterfly velocity that matches the results of the other two methods,
@@ 4.1 Background
In this section, we turn our attention to a bottom-up holographic EMD model and investigate the butterfly velocity at finite temperatures and chemical potentials in the dual QCD deconfined plasma. As mentioned in the Introduction, bottom-up models generally mimic QCD properties more accurately as compared to their higher-dimensional string-theory-inspired top-down counterparts; therefore, it is interesting to analyze the butterfly velocity in bottom-up models as well. To make our analysis more complete, we will consider another sophisticated bottom-up EMD-based QCD model in the next section.
Here we consider the bottom-up holographic QCD model of [53, 54]. The most general version of this model is a solution to the EMD gravity (3.1) containing Maxwell and dilaton matter fields. This model is based on the potential reconstruction technique [91, 92, 93, 94, 95, 96, 97, 98], where the potential is self-consistently fixed from field equations. In particular, for the following ansatz for the metric, dilaton, and gauge fields,
$ds^{2}=e^{2A(z)}\left(-z^{2}G(z)dt^{2}+\frac{dz^{2}}{z^{2}G(z)}+z^{2}(dx_{1}^{2}+dx^{2}+dx_{3}^{2})\right)\,$,
$\phi=\phi(z)\,,~~~\mathcal{A}_{M}=\Phi(z)\delta_{M}^{0}\,$, (4.1)
@@
Here $\mu$ is the chemical potential, and $a$ and $c$ are model parameters that are fixed by taking inputs from real QCD. In particular, by demanding the deconfinement transition to be around 0.270GeV0.270~GeV, as in the pure glue sector, fixes $a=0.15~\text{GeV}^{2}$, while the requirement of the linear Regge trajectory in the heavy-meson spectrum fixes $c=1.16~\text{GeV}^{2}$. Again, $z$ is the usual radial coordinate that runs from $z=z_{h}$ (horizon) to $z=\infty$ (boundary). The entropy and temperature of the black hole are given by
@@
There is another solution to this gravity model. This second solution corresponds to thermal-AdS, which does not have a horizon. The thermodynamic phase structure of the above solution reveals that there are two black hole solutions (large and small) that exist only above a minimum temperature $T_{min}$; below $T_{min}$, these black hole solutions cease to exist. The large/small black hole solution has positive/negative specific heat and is thermodynamically stable/unstable. Interestingly, there exists a Hawking/Page type phase transition between thermal-AdS and large black hole solutions as the temperature is varied. In particular, the large black hole phase is thermodynamically favored at high temperatures, whereas the thermal-AdS phase is favored at low temperatures. The corresponding phase transition temperature $T_{c}$ is a chemical potential-dependent quantity, and in particular, it decreases as the chemical potential increases. This is shown in Figure 5. In the dual boundary theory, the thermal-AdS and large black hole phases correspond to the confined and deconfined phases. Accordingly, this EMD gravity model provides a self-consistent bottom-up holographic model for the deconfinement transition, which mimics real QCD behavior quite well. For these reasons, along with their semianalytic nature, potential reconstruction-based holographic models have been widely employed to investigate various properties of QCD. These applications include studies of energy loss and drag forces [92], transport properties of QGP [101, 102], heavy quarkonium melting and dissociation [103, 104], chiral transition [105], etc. Since our main aim is to analyze the butterfly velocity and its thermal- and chemical-potential-dependent profile in the deconfined plasma phase, we will mainly concentrate on the large black hole solution in the next subsections.
Figure 5: Variation of deconfinement transition temperature $T_{c}$ with respect to $\mu$. In units of MeV.
It is also imperative to compare this holographic model with those of the 1RCBH and 2RCBH models considered earlier. From the QCD phenomenology point of view, the existence of the Hawking/Page phase transition allows us to consistently define the deconfinement transition temperature in the current model - a major improvement over the 1RCBH and 2RCBH models, where no such transition exists. Similarly, the dual confined phase, corresponding to the thermal-AdS background in the gravity side [obtained by taking the limit $z_{h}\rightarrow 0$ in Eq. (4.5)], exhibits not only the area law of the Wilson loop but also the linear Regge trajectory for the heavy-meson mass spectrum. The latter feature, in particular, is absent in 1RCBH and 2RCBH models. These are well-established properties of the QCD confined phase, thereby making the current holographic model highly suitable from the QCD point of view. Having said that, it is also important to keep in mind that the potential reconstruction-based holographic QCD models should be understood as approximate models, as the on-shell dilaton potential depends mildly on the model parameters.